A cosmic-ray proton in interstellar space has an energy of 23.0 MeV and executes a circular orbit having a radius equal to that of Mars' orbit around the Sun (2.28 1011 m). What is the magnetic field in that region of space? T
Answer #1

Radius, R = 2.28 x 10^11 m

Kinetic energy, KE = 23 MeV = 23 x 10^6 x 1.6 x 10^-19 = 3.68 x 10^-12 J

KE = 0.5 mv^2

Velocity, v = sqrt (2 x 3.68 x 10^-12/1.67 x 10^-27) = 6.64 x 10^7 m/s

Magnetic field, B = mv/qr

B = (1.67 x 10^-27 x 6.64 x 10^7)/(1.6 x 10^-19 x 2.28 x 10^11) = 3.04 x 10^-12 T

Comment in case any doubt please rate my answer....